1. A buffer solution is 0.309 M in H3PO4 and 0.241 M in KH2PO4. If Kal for H3PO4 is 7.5x10^-3, what is the pH of this buffer solution?

2. A 17.4 mL sample of a 0.308 M aqueous hydrocyanic acid solution is titrated with a 0.300 M aqueous barium hydroxide solution. What is the pH at the start of the titration, before any barium hydroxide has been added?

3. When a 15.1 mL sample of a 0.479 M aqueous nitrous acid solution is titrated with a 0.483 M aqueous sodium hydroxide solution, what is the pH after 22.5 mL of sodium hydroxide have been added?

4. A 24.1 mL sample of 0.286 M triethylamine, (C2H5)3N, is titrated with 0.218 M perchloric acid.

At the titration midpoint, the pH is

5. A 27.0 mL sample of 0.359 M trimethylamine, (CH3)3N, is titrated with 0.217 M hydroiodic acid.

After adding 63.0 mL of hydroiodic acid, the pH is
Answer #1

1)

Ka = 7.5*10^-3

pKa = - log (Ka)

= - log(7.5*10^-3)

= 2.125

use:

pH = pKa + log {[conjugate base]/[acid]}

pH = pKa + log {[H2PO4-]/[H3PO4]}

= 2.125+ log {0.241/0.309}

= 2.017

Answer: 2.02

Only 1 question at a time please